# Enhanced Generative Recommendation via Content and Collaboration Integration

### v1

@@ 3. Notations and Task Formulation
We first introduce the notations used in this paper. Then, we describe the task definition of generative recommendation.
@@ 3.2. Generative Recommendation
(1) $p(u,i)=\prod_{t=1}^{l}p(z_{i}^{t}|z_{i}^{1},z_{i}^{2},\cdots,z_{i}^{t-1},\mathcal{I}_{u}^{+}).$
@@ 4. Methodology
In this section, we describe the details of ColaRec. We first provide overview of the proposed ColaRec. Then the construction of items’ GIDs are detailed. After that, the user-item recommendation task and the item-item indexing task are described. Finally, we describe the joint optimization of above tasks.
@@ 4.1. Overview of ColaRec
Figure 3 illustrates an overview of the proposed ColaRec. ColaRec constructs the GID using a graph-based CF model, which effectively captures collaborative signals. The training of ColaRec consists of two tasks: the user-item recommendation task and the item-item indexing task. The user-item recommendation aims to map the content information of the user’s historical interacted items into the GID of the recommended item. The item-item indexing task targets on the mapping from the item side information into the item’s GID. Both the two tasks are achieved through a shared encoder-decoder based language model to better capture the textual content information. To this end, the recommendation task unifies both collaborative signals and item content information for better recommendation, while the indexing task performs the alignment between collaborative signals and content information. Note that parameters of the involved language model are also fine-tuned to better adapt the language model for recommendation.
@@ 4.2. Generative Identifier Construction
The construction of GIDs plays a crucial role for generative recommendation. Generally speaking, GIDs should satisfy the following expectations for better recommendation: (i) GIDs need to contain knowledge about both collaborative signals and content information; (ii) correlated items (e.g., similar items in content or items interacted by similar users) should have correlated GIDs; (iii) each item should have one unique GID and each GID should correspond to one specific item.
To fulfil above expectations, we utilize a hierarchical clustering approach to construct GIDs from a graph-based CF model. Specifically, we first extract item representations from a pretrained LightGCN model. Then the constrained $K$-means algorithm is called hierarchically based on the item representations. For the $t$-th level clustering with $t\in[1,l-1]$, the number of clusters is set to $K^{l-t}$. The next level clustering is conducted with items in the current cluster as the whole instance set. Regarding the last level of leaf node, we randomly allocate 1 to $K$ to the item. In this way, we establish a $K$-ary tree to organize the item set. Each item corresponds to a leaf node, while the path from the root to the leaf node is the GID of the item. Since LightGCN is trained on the user-item interaction graph, the GID can naturally encodes collaborative signals. Meanwhile, there is a codebook embedding matrix for every position of the GID, which will incorporate the content information in the item indexing task. We give detailed description in section 4.4. To this end, the GID together with the corresponding codebook embedding helps the recommender to model both collaborative signals and content information.
@@ 4.3. User-Item Recommendation
Model Inputs. The input sequence of user-item recommendation for user $u$ consists of unordered tuples with each tuple describing the content information of one interacted item of this user. For textual description of item $i$, we adopt the universal data format from (Li et al., 2023b). Specifically, textual description $c_{i}$ of item $i$ is formulated as an item “sentence” that comes from a flatted attribute dictionary consisting of key-value attribute pairs $(k,v)$, i.e., $[k_{1}{:}v_{1},k_{2}{:}v_{2},\dots]$. Besides, we also introduce the item atomic identifier $iad_{i}$ into the content information to further increase model fidelity. To this end, the content tuple of item $i$ is formulated as:
@@
Since the training of ColaRec has two tasks, in the user-item recommendation task, we augment the input by adding a textual “user prompt” (denoted as $prompt_{u}$) at the beginning of the input to inform the model that the ongoing task is the recommendation task. Thus, the input for the user-item recommendation task is:
(3) $X_{u}=[prompt_{u},\{c_{i}|i\in\mathcal{I}_{u}^{+}\}].$
Language Models for Recommendation. We employ an encoder-decoder based language model to capture textual content information. Given the model input $X_{u}$, the language model encoder captures the semantic information of $X_{u}$ and returns the hidden state $\text{Encoder}(X_{u})$. After that, given the generated tokens $z^{<t}$ before the $t$-th generation step, the decoder generates the latent representation $\mathbf{d}_{t}\in\mathbb{R}^{m}$ for the $t$-th token of GID. $m$ is the dimension of the latent representation. This process can be formulated as:
@@
In this work, we use a pretrained T5 (Raffel et al., 2020) as the language model. The parameters of T5 are also fine-tuned through back propagation to better adapt the language model for recommendation.
@@ 4.4. Item-Item Indexing
Model Inputs. The input sequence for item indexing contains the textual information of the item. Besides, we also introduce information of users who have interacted with this item, to further encode collaborative signals. Similar to the recommendation task, we augment the input an “item prompt” (denoted as $prompt_{i}$) at the beginning of the input. Therefore, the input of the indexing task is formulated as:
(7) $X_{i}=[prompt_{i},c_{i},\{uad_{u}|u\in\mathcal{U}_{i}^{+}\}].$
Language Models for Indexing. The indexing task is conducted through the same language model and codebook embeddings as the recommendation task. The generation probability for the indexing task is formulated similarly with Eq. (4) and Eq. (5) except that the model input is $X_{i}$ instead of $X_{u}$. We adopt the cross-entropy loss for parameter tuning. The loss for item indexing is defined as:
@@
Contrastive Optimization. To conduct better alignment between collaborative signals and content information, a contrastive loss is further introduced. The idea is that items with similar GIDs should also be similar in the content-based semantic space. To this end, for the item $i$, we randomly sample an item $i_{+}$ which has overlapped sub-sequence in GIDs as the positive sample, and another randomly sampled item $i_{-}$ without overlapped GID tokens as the negative sample. The contrastive loss is defined as:
@@
where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$), and $\sigma$ denotes the sigmoid function. Such a contrastive loss helps the encoder to learn better item input representations.
4.5. Joint Optimization
Besides the above tasks, we further introduce a ranking loss to enhance the ranking ability of ColaRec. For a $(u,i)$ pair in the training dataset, we sample one item that the user $u$ has not interacted with, and for which the GID has no overlap with the positive item $i$, as the negative sample $i_{-}$. The BPR loss (Rendle et al., 2012) is utilized to optimize the ranking, which is formulated as:
(10) $\mathcal{L}_{\text{BPR}}=-\ln\sigma(\mathbf{h}(X_{u})\cdot(\mathbf{h}(X_{i})-\mathbf{h}(X_{i_{-}}))).$
where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$). The above loss pushes together the positive $(u,i)$ pair in the semantic space and pushes away the negative $(u,i_{-})$ pair, thus helping the encoder to capture the ranking knowledge between items.
Finally, ColaRec is trained with the above-described tasks jointly:
@@ 4.5. Joint Optimization
During the inference, recommendation lists are generated via beam search. To avoid the recommender from generating invalid GIDs that cannot be mapped to concrete items in the candidate set, we employ the constrained beam search (Cao et al., 2021) to limit the generative range of the current token based on the prefix tokens.
Table 1. Statistics of three public datasets after preprocessing.
@@
Table 2. Performance comparison on three public datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. * denotes the paired t-test with significance p-value ¡ 0.1.
CF-based Methods Generative Methods
Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER Ours
Beauty Recall@5 0.0447 0.0649 0.0551 0.0650 0.0530 0.0524 0.0128 0.0451 0.0519 0.0667*
Recall@10 0.0653 0.0952 0.0831 0.0940 0.0776 0.0741 0.0228 0.0705 0.0799 0.0993*
Recall@20 0.0889 0.1314 0.1193 0.1320 0.1093 0.1016 0.0360 0.1018 0.1154 0.1371*
NDCG@5 0.0315 0.0450 0.0377 0.0452 0.0362 0.0378 0.0084 0.0305 0.0350 0.0449
NDCG@10 0.0383 0.0549 0.0469 0.0547 0.0443 0.0450 0.0117 0.0385 0.0443 0.0556*
NDCG@20 0.0445 0.0643 0.0563 0.0646 0.0526 0.0521 0.0151 0.0470 0.0534 0.0654*
Sports Recall@5 0.0206 0.0418 0.0355 0.0427 0.0314 0.0273 0.0117 0.0320 0.0374 0.0442*
Recall@10 0.0321 0.0623 0.0557 0.0631 0.0476 0.0403 0.0178 0.0497 0.0572 0.0660*
Recall@20 0.0471 0.0901 0.0836 0.0908 0.0713 0.0569 0.0284 0.0766 0.0881 0.0964*
NDCG@5 0.0140 0.0288 0.0240 0.0294 0.0208 0.0193 0.0079 0.0225 0.0249 0.0294
NDCG@10 0.0177 0.0355 0.0306 0.0361 0.0261 0.0235 0.0099 0.0284 0.0313 0.0364
NDCG@20 0.0215 0.0426 0.0377 0.0431 0.0321 0.0278 0.0126 0.0350 0.0392 0.0442*
Phone Recall@5 0.0410 0.0713 0.0643 0.0717 0.0569 0.0470 0.0187 0.0412 0.0601 0.0745*
Recall@10 0.0603 0.1052 0.0976 0.1043 0.0855 0.0668 0.0341 0.0625 0.0895 0.1121*
Recall@20 0.0871 0.1487 0.1420 0.1481 0.1233 0.0928 0.0564 0.0966 0.1299 0.1587*
NDCG@5 0.0282 0.0481 0.0423 0.0488 0.0378 0.0315 0.0121 0.0282 0.0403 0.0490
NDCG@10 0.0344 0.0590 0.0530 0.0593 0.0470 0.0379 0.0170 0.0347 0.0498 0.0611*
NDCG@20 0.0412 0.0700 0.0643 0.0704 0.0566 0.0445 0.0225 0.0431 0.0600 0.0729*
5. Experiment Settings
In this section, we describe experiment settings to evaluate the proposed ColaRec. We aim to answer the following research questions:
@@ 5.1. Datasets
We use three real-world public datasets from Amazon Product Reviews [^3] https://jmcauley.ucsd.edu/data/amazon/, one of the most widely used recommendation benchmarks, to evaluate the performance of ColaRec. In particular, the experiments are conducted on three subcategories, including “Beauty”, “Sports and Outdoors”, and “Cell Phones and Accessories”. Users and items which have less than five interactions are filtered out. Table 1 shows the statistics of all three datasets after preprocessing. As for content information, we select “title”, “brand” and “categories” from the item metadata as the textual content description of items.
@@ 5.2. Evaluation Protocols
We adopt cross-validation to evaluate the performance of recommnders. In this paper we target on general recommendation other than sequential recommendation. To this end, we randomly split each user’s historical interactions into the training/validation/test set with the ratio of 8:1:1. We employ two widely used metrics, recall@$n$ and normalized discount cumulative gain (NDCG@$n$), to evaluate the model performance. Recall evaluates how many ground-truth items occur in the recommended list, while NDCG further focuses on their rankings in the list. Note that in this paper, the candidate item set is the whole item set, other than a small subset with selected items. Each experiment is conducted three times and the average score is reported.
@@ 5.3. Baselines
We compare ColaRec with several representative related baselines, including both conventional CF-based methods and generative methods. CF-based baselines include NeuMF (He et al., 2017), LightGCN (He et al., 2020), SimpleX (Mao et al., 2021), and NCL (Lin et al., 2022). Generative baselines include MultiVAE (Liang et al., 2018), DiffRec (Wang et al., 2023b), DSI (Tay et al., 2022), and TIGER (Rajput et al., 2023).
Table 3. Performance comparison of long-tail users on three public datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. ** denotes that the improvements are significant with p-value ¡ 0.05.
CF-based Methods Generative Methods
Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER Ours
Beauty Recall@5 0.0416 0.0636 0.0555 0.0639 0.0510 0.0464 0.0131 0.0415 0.0519 0.0660**
Recall@10 0.0604 0.0922 0.0825 0.0907 0.0742 0.0662 0.0228 0.0653 0.0799 0.0975**
Recall@20 0.0817 0.1253 0.1160 0.1264 0.1039 0.0917 0.0354 0.0940 0.1154 0.1327**
Sports Recall@5 0.0209 0.0433 0.0355 0.0440 0.0329 0.0267 0.0116 0.0307 0.0380 0.0456**
Recall@10 0.0317 0.0639 0.0562 0.0645 0.0495 0.0394 0.0170 0.0472 0.0581 0.0674**
Recall@20 0.0468 0.0904 0.0836 0.0908 0.0725 0.0553 0.0273 0.0728 0.0882 0.0976**
Phone Recall@5 0.0405 0.0723 0.0660 0.0727 0.0571 0.0451 0.0206 0.0404 0.0602 0.0756**
Recall@10 0.0590 0.1054 0.0986 0.1043 0.0861 0.0641 0.0371 0.0623 0.0898 0.1131**
Recall@20 0.0855 0.1482 0.1418 0.1473 0.1228 0.0899 0.0600 0.0939 0.1293 0.1590**
@@
  • DiffRec (Wang et al., 2023b) is a new proposed recommendation model based on diffusion models. DiffRec learns the user-item interaction knowledge through a recontruction and denoising manner.
  • DSI (Tay et al., 2022) is a generative document retrieval method. To adapt DSI for recommendation, we formulate the input as GIDs of the user’s historical interacted items. We use two versions of DSI. DSI-R refers to the DSI model with a random string as the GID of the item. DSI-S is a DSI model which constructs the item GID with a hierarchical $K$-means algorithm based on the item textual content embeddings from a pre-trained BERT model.
  • TIGER (Rajput et al., 2023) is a generative recommendation method. Specifically, a pre-trained Sentence-T5 encoder is used to obtain embeddings of the item’s textual content. These embeddings are then quantized using an RQ-VAE to build GIDs. We do not introduce sequential orders to adapt TIGER for general recommendation.

### v2

@@ 3.2. Generative Recommendation
(1) $p(u,i)=\prod_{t=1}^{l}p(z_{i}^{t}|\mathcal{I}_{u}^{+},z_{i}^{1},z_{i}^{2},\cdots,z_{i}^{t-1}).$
@@ 4. Methodology
In this section, we describe the details of ColaRec. We first provide the overview of the proposed ColaRec. Then the construction of items’ GIDs is detailed. After that, the user-item recommendation task and the item-item indexing task are described. Finally, we describe the joint optimization of the above tasks.
@@ 4.1. Overview of ColaRec
Figure 3 illustrates an overview of the proposed ColaRec. ColaRec constructs the GID using a graph-based CF model, which effectively captures collaborative signals. The training of ColaRec consists of two tasks: the user-item recommendation task and the item-item indexing task. The user-item recommendation aims to map the content information of the user’s historical interacted items into the GID of the recommended item. The item-item indexing task targets on the mapping from the item side information into the item’s GID. Both the two tasks are achieved through a shared encoder-decoder based model. The recommendation task unifies both collaborative signals and item content information for better recommendation, while the indexing task performs the alignment between collaborative signals and content information.
@@ 4.2. Generative Identifier Construction
The construction of GIDs plays a crucial role in generative recommendation. Generally speaking, GIDs should satisfy the following expectations for better recommendation: (i) GIDs need to contain knowledge about both collaborative signals and content information; (ii) Correlated items (e.g., similar items in content or items interacted by similar users) should have correlated GIDs; (iii) Each item should have one unique GID and each GID should correspond to one specific item.
To fulfill above expectations, we utilize a hierarchical clustering approach to construct GIDs from a graph-based CF model. Specifically, we first extract item representations from a pretrained LightGCN model. Then the constrained $K$-means algorithm is called hierarchically based on the item representations. The next level clustering is conducted with items in the current cluster as the whole item set. For the $t$-th level clustering with $t\in[1,l-1]$, the number of items is no more than $K^{l-t}$ in each cluster. Regarding the last level of leaf nodes, we randomly allocate 1 to $K$ to the items. In this way, we establish a $K$-ary tree to organize the item set. Each item corresponds to a leaf node, while the path from the root to the leaf node is the GID of the item. Since LightGCN is trained on the user-item interaction graph, the GID can naturally encode collaborative signals. Meanwhile, there is a codebook embedding matrix for every position of the GID, which will incorporate the content information in the item indexing task. We give the detailed description in section 4.4. To this end, the GID together with the corresponding codebook embeddings helps the recommender to model both collaborative signals and content information.
@@ 4.3. User-Item Recommendation
Model Inputs. The input sequence of user-item recommendation for user $u$ consists of unordered tuples with each tuple describing the content information of one interacted item of this user. For the textual description of item $i$, we adopt the universal data format from (Li et al., 2023b). Specifically, textual description $c_{i}$ of item $i$ is formulated as a sequence that comes from a flatted attribute dictionary consisting of key-value attribute pairs $(k,v)$, i.e., $[k_{1}{:}v_{1},k_{2}{:}v_{2},\dots]$. Besides, we also introduce the item atomic identifier $iad_{i}$ into the content information to further increase model fidelity. To this end, the content tuple of item $i$ is formulated as:
@@
Since the training of ColaRec has two tasks, in the user-item recommendation task, we augment the input with a special task token $task_{u}$ at the beginning of the input to inform the model that the ongoing task is the recommendation task. Thus, the input for the user-item recommendation task is:
(3) $X_{u}=[task_{u},\{c_{i}|i\in\mathcal{I}_{u}^{+}\}].$
Item Generation. We employ an encoder-decoder based Transformer model to generate item recommendation. Given the model input $X_{u}$, the model encoder captures the semantic information of $X_{u}$ and returns the hidden state $\text{Encoder}(X_{u})$. After that, given the generated tokens $z^{<t}$ before the $t$-th generation step, the decoder generates the latent representation $\mathbf{d}_{t}\in\mathbb{R}^{m}$ for the $t$-th token of GID. $m$ is the dimension of the latent representation. This process can be formulated as:
@@
In this work, we use a pretrained T5 (Raffel et al., 2020) as the Transformer model. The parameters of T5 are also fine-tuned through back propagation to better adapt the model for recommendation. Note that our approach is model-agnostic and T5 can be replaced by other sequence-to-sequence models, even without pretraining.
@@ 4.4. Item-Item Indexing
Model Inputs. The input sequence for item indexing contains the textual information of the item. Besides, we also introduce information of users who have interacted with this item, to further encode collaborative signals. Similar to the recommendation task, we augment the input with a special task token $task_{i}$ at the beginning of the input for item indexing. Therefore, the input of item indexing is formulated as:
(7) $X_{i}=[task_{i},c_{i},\{uad_{u}|u\in\mathcal{U}_{i}^{+}\}].$
Item Indexing. The indexing task is conducted through the same model and codebook embeddings as the recommendation task. The generation probability for the indexing task is formulated similarly with Eq. (4) and Eq. (5) except that the model input is $X_{i}$ instead of $X_{u}$. We adopt the cross-entropy loss for parameter learning. The loss for item indexing is defined as:
@@
4.5. Multi-Task Training
Besides the above two tasks, we further introduce a ranking loss to enhance the ranking ability of ColaRec and a contrastive loss to conduct better alignment.
Item Ranking. For a $(u,i)$ pair in the training dataset, we randomly sample one item that the user $u$ has not interacted with as the negative sample $i_{-}$. The BPR loss (Rendle et al., 2012) is utilized to optimize the item ranking, which is formulated as:
(9) $\mathcal{L}_{\text{bpr}}=-\ln\sigma(\mathbf{h}(X_{u})\cdot(\mathbf{h}(X_{i})-\mathbf{h}(X_{i_{-}}))),$
where $\mathbf{h}(\cdot)$ denotes the last hidden states of Encoder($\cdot$), and $\sigma$ denotes the sigmoid function. The above loss pushes together the positive $(u,i)$ pair and pushes away the negative $(u,i_{-})$ pair, thus helping the model to capture the ranking knowledge between items.
Contrastive Learning. To conduct better alignment between collaborative signals and content information, a contrastive loss is further introduced. The idea is that items with similar GIDs should also be similar in the content-based semantic space. To this end, for the item $i$, we randomly sample an item $i_{+}$ which has the overlapped prefix tokens in GIDs as the positive sample. The sampled item $i_{-}$ in $\mathcal{L}_{\text{bpr}}$ serves as the negative sample. Note that here we ensure the sampled item $i_{-}$ in $\mathcal{L}_{\text{bpr}}$ has no overlapped GID tokens with item $i$. The contrastive loss is defined as:
@@ 4.5. Multi-Task Training
Such a contrastive loss helps the model to learn better item input representations.
Joint Optimization. Finally, ColaRec is trained with the above described tasks jointly:
@@
During the inference, to avoid the recommender from generating invalid GIDs, we employ the constrained beam search (Cao et al., 2021) to limit the generative range of the current token based on the prefix tokens.
Table 1. Statistics of four public datasets after preprocessing.
@@
Recipe 17,813 41,240 555,618
Table 2. Performance comparison on four public datasets. The best and the second-best scores are marked in bold and underlined fonts, respectively. * denotes the paired t-test with significance p-value ¡ 0.1. R and N stand for Recall and NDCG.
CF-based Methods Generative Models for Recommendation
Datasets Metric NeuMF LightGCN SimpleX NCL MutiVAE DiffRec DSI-R DSI-S TIGER LC-Rec Ours
Beauty R@5 0.0447 0.0649 0.0551 0.0650 0.0530 0.0524 0.0128 0.0451 0.0519 0.0492 0.0667*
R@10 0.0653 0.0952 0.0831 0.0940 0.0776 0.0741 0.0228 0.0705 0.0799 0.0770 0.0993*
R@20 0.0889 0.1314 0.1193 0.1320 0.1093 0.1016 0.0360 0.1018 0.1154 0.1104 0.1371*
N@5 0.0315 0.0450 0.0377 0.0452 0.0362 0.0378 0.0084 0.0305 0.0350 0.0326 0.0449
N@10 0.0383 0.0549 0.0469 0.0547 0.0443 0.0450 0.0117 0.0385 0.0443 0.0415 0.0556*
N@20 0.0445 0.0643 0.0563 0.0646 0.0526 0.0521 0.0151 0.0470 0.0534 0.0499 0.0654*
Sports R@5 0.0206 0.0418 0.0355 0.0427 0.0314 0.0273 0.0117 0.0320 0.0374 0.0397 0.0442*
R@10 0.0321 0.0623 0.0557 0.0631 0.0476 0.0403 0.0178 0.0497 0.0572 0.0617 0.0660*
R@20 0.0471 0.0901 0.0836 0.0908 0.0713 0.0569 0.0284 0.0766 0.0881 0.0931 0.0964*
N@5 0.0140 0.0288 0.0240 0.0294 0.0208 0.0193 0.0079 0.0225 0.0249 0.0264 0.0294
N@10 0.0177 0.0355 0.0306 0.0359 0.0261 0.0235 0.0099 0.0284 0.0313 0.0335 0.0364*
N@20 0.0215 0.0426 0.0377 0.0431 0.0321 0.0278 0.0126 0.0350 0.0392 0.0413 0.0442*
Phone R@5 0.0410 0.0713 0.0643 0.0717 0.0569 0.0470 0.0187 0.0412 0.0601 0.0615 0.0745*
R@10 0.0603 0.1052 0.0976 0.1043 0.0855 0.0668 0.0341 0.0625 0.0895 0.0919 0.1121*
R@20 0.0871 0.1487 0.1420 0.1481 0.1233 0.0928 0.0564 0.0966 0.1299 0.1354 0.1587*
N@5 0.0282 0.0481 0.0423 0.0486 0.0378 0.0315 0.0121 0.0282 0.0403 0.0408 0.0490*
N@10 0.0344 0.0590 0.0530 0.0593 0.0470 0.0379 0.0170 0.0347 0.0498 0.0506 0.0611*
N@20 0.0412 0.0700 0.0643 0.0704 0.0566 0.0445 0.0225 0.0431 0.0600 0.0615 0.0729*
Recipe R@5 0.0118 0.0188 0.0114 0.0192 0.0167 0.0142 0.0142 0.0157 0.0168 0.0174 0.0198*
R@10 0.0210 0.0296 0.0202 0.0298 0.0285 0.0235 0.0248 0.0270 0.0292 0.0289 0.0306*
R@20 0.0339 0.0454 0.0328 0.0459 0.0462 0.0343 0.0403 0.0436 0.0464 0.0454 0.0482*
N@5 0.0088 0.0149 0.0093 0.0149 0.0128 0.0105 0.0107 0.0122 0.0137 0.0138 0.0151*
N@10 0.0119 0.0182 0.0122 0.0182 0.0167 0.0135 0.0141 0.0158 0.0176 0.0175 0.0185*
N@20 0.0154 0.0223 0.0156 0.0224 0.0214 0.0165 0.0182 0.0202 0.0221 0.0218 0.0232*
5. Experiment
In this section, we conduct experiments to evaluate the proposed ColaRec. We aim to answer the following research questions:
@@ 5.1. Datasets
We use four real-world public datasets to evaluate the performance of ColaRec. Specifically, the experiments are conducted on three subcategories from Amazon Product Reviews [^4] https://jmcauley.ucsd.edu/data/amazon/ (“Beauty”, “Sports and Outdoors”, and “Cell Phones and Accessories”) and “Recipe” from Food.com [^5] https://www.kaggle.com/datasets/shuyangli94/food-com-recipes-and-user-interactions. Users and items that have less than five interactions are filtered out. Table 1 shows the statistics of all four datasets. As for content information, we use “title”, “brand” and “categories” from the Amazon item metadata as the textual content information of items. For Recipe, we use “name”, “description”, and “tag” to describe item content.
@@ 5.2. Evaluation Protocols
We adopt cross-validation to evaluate the performance of recommenders. In this paper we focus on general recommendation rather than sequential recommendation. To this end, we randomly split each user’s historical interactions into the training/validation/test set with the ratio of 8:1:1. We employ two widely used metrics, recall@$n$ and normalized discount cumulative gain (NDCG@$n$), to evaluate the model performance. Recall evaluates how many ground-truth items occur in the recommended list, while NDCG further focuses on their rankings in the list. Note that in this paper, the candidate item set is the whole item set, other than a small subset with selected items. Each experiment is conducted three times and the average score is reported.
@@ 5.3. Baselines
We compare ColaRec with several representative related baselines, including both conventional CF-based methods and generative models for recommendation. CF-based baselines include NeuMF (He et al., 2017), LightGCN (He et al., 2020), SimpleX (Mao et al., 2021), and NCL (Lin et al., 2022). Generative models for recommendation include MultiVAE (Liang et al., 2018), DiffRec (Wang et al., 2023b), DSI (Tay et al., 2022), TIGER (Rajput et al., 2023) and LC-Rec(Zheng et al., 2023).
@@
  • DiffRec (Wang et al., 2023b) is a new proposed recommendation model based on diffusion models. DiffRec learns the user-item interaction knowledge through a reconstruction and denoising manner.
  • DSI (Tay et al., 2022) is a generative document retrieval method. To adapt DSI for recommendation, we formulate the input as GIDs of the user’s historical interacted items. We use two versions of DSI. DSI-R refers to the DSI model with a random string as the GID of the item. DSI-S is a DSI model which constructs the item GIDs with a hierarchical $K$-means algorithm based on the item textual content embeddings from a pretrained BERT model.
  • TIGER (Rajput et al., 2023) is a generative recommendation method. Specifically, a pretrained Sentence-T5 encoder is used to obtain embeddings of the items’ textual content. These embeddings are then quantized using an RQ-VAE to build GIDs. We do not introduce sequential orders to adapt TIGER for general recommendation.
  • LC-Rec (Zheng et al., 2023) follows a similar approach to TIGER (Rajput et al., 2023) for item index learning. Besides, it employs various language generation tasks under different prompts to accomplish recommendation.